Find the standard form of the equation of the line with slope 7 passing through the points (3, -7)
Answer by rosey51235(30)

If you want it in the formula y=mx+b, all you have to do is plug in what you're given already and solve for what's left.
x and y are given as the point it passes through (x,y) and m=slope (7) so just plug in and solve step by step!
-7=7(3)+b
-7=21+b
(-21 on both sides)
-28=b
So your final equation should be Y=7x-28
Hope I helped! :)
-Rosey